OpenStudy (ibbutibbu.):

Can someone help me with factoring? 4x^3 + 43x^2 + 30x

OpenStudy (solomonzelman):

first factor out of x, can you do that for me?

OpenStudy (ibbutibbu.):

x ( 4x^2 + 43x + 30 ) ?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now, lets focus on just the 4x²+43x+30

OpenStudy (solomonzelman):

you know that if you choose (2x )(2x ) that is not going to give an odd coefficient 43 in a middle term, therefore you have to choose (4x )(x )

OpenStudy (solomonzelman):

and you know that you have (4x + 1)(x + ) (2 pluses, because your trinomial is positive entirely)

OpenStudy (solomonzelman):

now you have to choose two constant factors of 30 to fill it in.

OpenStudy (ibbutibbu.):

The factors are 1, 2, 3, 5, 6, 10, 15, 30

OpenStudy (ibbutibbu.):

I think

OpenStudy (solomonzelman):

yes, possible picks 1 & 30 2 & 15 3 & 10 5 & 6

OpenStudy (solomonzelman):

4x²+43x+30 (4x + )(x + )

OpenStudy (ibbutibbu.):

The rule it that a+b=43 and ab=30?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

beautiful

OpenStudy (ibbutibbu.):

im not entirely sure

OpenStudy (solomonzelman):

30×1=30, but 30 & 1 don't make the a+b=43

OpenStudy (solomonzelman):

one note/addition

OpenStudy (solomonzelman):

`(4)a+(1)b=43` and `ab=30`

OpenStudy (ibbutibbu.):

the 10 +3

OpenStudy (solomonzelman):

you multiply on of the numbers by 4 and the other by 1 (multiplying by coefficients of x, just as you would if you were to expand it)

OpenStudy (ibbutibbu.):

bc 4(10) + 3 = 43

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

4x²+43x+30 (4x + )(x + ) fill it for me please...

OpenStudy (ibbutibbu.):

(4x + 10)(x + 3) ?

OpenStudy (solomonzelman):

wait, this would give a middle term of (10 • x) + (4x • 3) = 10x+12x=22x

OpenStudy (solomonzelman):

the other way around

OpenStudy (ibbutibbu.):

so (4x + 3)(x + 10) ?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

(4x•10)+(x•3)=40x+3x=43x

OpenStudy (solomonzelman):

4x³ + 43x² + 30x x ( 4x² + 43x + 30 ) x (4x + 3)(x + 10)

OpenStudy (solomonzelman):

` x (4x + 3)(x + 10) ` is it.... you can't factor it more. DONE

OpenStudy (ibbutibbu.):

thx so much

OpenStudy (solomonzelman):

yw